OpenStudy (anonymous):

find lim theta->pi/2 of cos theta/sin 2 theta

OpenStudy (anonymous):

\[\lim_{\theta \rightarrow \frac{ \pi }{ 2 }}\frac{ \cos \theta }{ \sin 2 \theta }\]

OpenStudy (anonymous):

\[\frac{ \cos \theta }{ 2\sin \theta \cos \theta }\]

OpenStudy (anonymous):

very good..

OpenStudy (anonymous):

\[\frac{ 1 }{ 2 \sin \theta }\]

OpenStudy (anonymous):

that's it .. you done :)

OpenStudy (anonymous):

now plug it into theta

OpenStudy (anonymous):

cant be as this gives me a division by 0.

OpenStudy (anonymous):

no.. sin(pi/2) = 1

OpenStudy (anonymous):

im a chop lol

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

thanks tho i just needed confimation

OpenStudy (anonymous):

yes youve done this by yourself

OpenStudy (anonymous):

lol yup i do this alot, caus then if i do make a mistake it can get caught and i dont teach myself rubbish

OpenStudy (anonymous):

How about a medal for @Razzputin everyone?

OpenStudy (anonymous):

ha ha thanks rajathsbhat tho i dnt see y im the one who needs the medal?

OpenStudy (anonymous):

you answered the question by yourself, you know..

OpenStudy (anonymous):

thanks anyway hey. i g2g work to do... have fun ppl thanks for all the help